# Flowception: Temporally Expansive Flow Matching for Video Generation

### v1

@@ 4.3 Ablation experiments
Impact of learned insertions.  To assess the importance of the insertion head that learns to predict where to insert frames, we compare it to inserting frames in a data-independent manner. In particular, we first determine the video length $n$. We then consider inserting frames in a random order, using a “hierarchical” scheme where we iteratively insert them in the middle of the largest interval in $\{1,\dots,n\}$ where no frame has been inserted yet, or using a left-to-right pattern as used in AR models. We present the results of this experiment in Table˜2. We find that across metrics, results improve as we move from the random pattern to the hierarchical pattern, to the left-to-right, and finally to the data-driven insertion patterns in Flowception .
Note that for the left-to-right pattern, the FVD results in Table˜2 (23.61) are much better than the FVD for the AR baselines in Table˜4 (47.48 and 45.13). We hypothesize that this difference is due to the fact that the AR baselines denoise the next frame once previous ones are already fully denoised and committed to, while in Flowception new frames start denoising while previous ones are still not fully denoised (more similar to the full-sequence baseline which yields a FVD of 26.17 on RealEstate10K).
Insertion guidance.  Classifier-free guidance (CFG) is commonly used to achieve better prompt alignment and image quality in diffusion and flow models Ho and Salimans (2021). Similarly, in Flowception , we can perform CFG on the insertion rates. Following CFG, we randomly drop the conditional information $c$ during training, and define the guided update as
@@
where $w_{s}\geq 1$ is the guidance scale for the insertion rate. Figure˜7 presents a scatter plot comparing video lengths for I2V obtained with and without guidance, initiating generation from the same seed for each starting frame. For efficiency we limit the number of frames to 20 in this experiment. The result clearly indicates that using rate guidance ($w_{s}\!=\!5$) biases the model towards generating longer videos.
@@
Without guidance, or using lower values, we find insertion guidance to prevent problems of under-insertions which can result in a choppy transition between two frames. This is confirmed in Table˜3 where motion smoothness increases while dynamic degree decreases as $w_{s}$ increases.
@@
Causality in AR models.  In the literature, most autoregressive video generation models make use of frame-wise causal attention Deng et al. (2024); Chen et al. (2025), which allows for KV caching and therefore efficient inference for autoregressive generation. Other works improve on quality and long term rollout consistency by using bidirectional attention in the context and applying frame-wise noising for each frame Song et al. (2025). Such an approach, while effective, significantly raises inference time expected compute from $O(L^{2}n^{2})$ to $O(L^{2}n^{3})$ where $n$ is the number of video frames, and $L$ is the number of tokens per frame. A more detailed analysis of FLOPs is provided in the supplementary material.
We provide a comparison between both autoregressive methods in Table˜4. The results demonstrate slightly better performance of the non-causal attention pattern, outperforming the causal one on every metric. Flowception yields further improvements over the non-causal AR baseline on all metrics but subject consistency.
Table 4: Comparison of autoregressive attention patterns. Autoregressive baseline with differnt attention amskes and comparison to Flowception in terms of VBench metrics and FVD, Models trained on Real Estate10K.
@@
Local vs. global attention.  Flowception generates a video by progressively inserting new frames. Empirically, we observe that Flowception tends to insert distant frames first, and then progressively “fills in the blanks”. This suggests that Flowception could be more amenable to using efficient local attention patterns, as far-away frames can still attend to each other early in the denoising process when the sequence is still short. To assess whether Flowception is more amenable to using local attention windows than the full-sequence baseline, we compare models using global attention and using local attention windows restricted to the previous and next $K$ frames in the sequence. The results in Figure˜8 indicate that Flowception suffers far less from using local attention windows than the full-sequence model, presumably due the communication between distant frames early in the Flowception denoising process, when the intermediate frames have not yet been inserted, even when using small attention windows.

### v2

@@ 4.3 Ablation experiments
Impact of learned insertions.  To assess the importance of the insertion head that learns to predict where to insert frames, we compare it to inserting frames in a data-independent manner. In particular, we first determine the video length $n$. We then consider inserting frames in a random order, using a “hierarchical” scheme where we iteratively insert them in the middle of the largest interval in $\{1,\dots,n\}$ where no frame has been inserted yet, or using a left-to-right pattern as used in AR models. We present the results of this experiment in Table˜3. We find that across metrics, results improve as we move from the random pattern to the hierarchical pattern, to the left-to-right, and finally to the data-driven insertion patterns in Flowception.
Note that for the left-to-right pattern, the FVD results in Table˜3 (23.61) are much better than the FVD for the AR baselines in Table˜5 (47.48 and 45.13). We hypothesize that this difference is due to the fact that the AR baselines denoise the next frame once previous ones are already fully denoised and committed to, while in Flowception new frames start denoising while previous ones are still not fully denoised (more similar to the full-sequence baseline which yields a FVD of 26.17 on RealEstate10K).
Insertion guidance.  Classifier-free guidance (CFG) is commonly used to achieve better prompt alignment and image quality in diffusion and flow models (Ho and Salimans, 2021). Similarly, in Flowception, we can perform CFG on the insertion rates. Following CFG, we randomly drop the conditional information $c$ during training, and define the guided update as
@@
where $w_{s}\geq 1$ is the guidance scale for the insertion rate. Figure˜8 presents a scatter plot comparing video lengths for I2V obtained with and without guidance, initiating generation from the same seed for each starting frame. For efficiency we limit the number of frames to 20 in this experiment. The result clearly indicates that using rate guidance ($w_{s}\!=\!5$) biases the model towards generating longer videos.
@@
Without guidance, or using lower values, we find insertion guidance to prevent problems of under-insertions which can result in a choppy transition between two frames. This is confirmed in Table˜4 where motion smoothness increases while dynamic degree decreases as $w_{s}$ increases.
@@
Causality in AR models.  In the literature, most autoregressive video generation models make use of frame-wise causal attention (Deng et al., 2025; Chen et al., 2025), which allows for KV caching and therefore efficient inference for autoregressive generation. Other works improve on quality and long term rollout consistency by using bidirectional attention in the context and applying frame-wise noising for each frame (Song et al., 2025). Such an approach, while effective, significantly raises inference time expected compute from $O(L^{2}n^{2})$ to $O(L^{2}n^{3})$ where $n$ is the number of video frames, and $L$ is the number of tokens per frame. A more detailed analysis of FLOPs is provided in the supplementary material.
We provide a comparison between both autoregressive methods in Table˜5. The results demonstrate slightly better performance of the non-causal attention pattern, outperforming the causal one on every metric. Flowception yields further improvements over the non-causal AR baseline on all metrics but subject consistency.
Table 5: Comparison of autoregressive attention patterns. Autoregressive baseline with different attention masks and comparison to Flowception in terms of VBench metrics and FVD. Models trained on RealEstate10K.
@@
Local vs. global attention.  Flowception generates a video by progressively inserting new frames. Empirically, we observe that Flowception tends to insert distant frames first, and then progressively “fills in the blanks”. This suggests that Flowception could be more amenable to using efficient local attention patterns, as far-away frames can still attend to each other early in the denoising process when the sequence is still short. To assess whether Flowception is more amenable to using local attention windows than the full-sequence baseline, we compare models using global attention and using local attention windows restricted to the previous and next $K$ frames in the sequence. The results in Figure˜9 indicate that Flowception suffers far less from using local attention windows than the full-sequence model, presumably due the communication between distant frames early in the Flowception denoising process, when the intermediate frames have not yet been inserted, even when using small attention windows.
@@ 5 Conclusion
Acknowledgements.  Karteek Alahari was supported in part by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korean Government (MSIT) (No. RS-2024-00457882, National AI Research Lab Project).